Skip to content

Worklog

Current Status

| Last updated | by | Status | Reason | |--------------|----------------------------------------|-----------------------------------------------------------------|----------------------------------------------------| | 2016-06-10 | [Greg](User:gregory.morris "wikilink") | Partially Working | High acceleration on Y-axis can cause missed steps |
Work Log -------- | When | Work Done | Who By | |------------|----------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------|----------------------------------------| | 2015-09-21 | Reconnected up/down switch for moving bed. | [Bob](User:parag0n "wikilink") | | 2015-09-23 | Added lid microswitch to turn off laser when lid is opened. | [Bob](User:parag0n "wikilink") | | 2015-11-15 | Added reset microswitch. Wired coolant system into network. Replaced stepper power supply with 30V one. Blew up one of the stepper drivers. Fuck. Wired up routoutcnc stepper driver to replace broken driver, immediately blew up too. Fuck. Wired up second routoutcnc stepper driver to 12v instead of 30V. MBed is buggered, heating up and crashing, but cutter works for a short time. | [Bob](User:parag0n "wikilink") | | 2015-11-20 | Replaced MBed with new one, replaced other stepper driver with third (and last!) routoutcnc board on 12V. Transferred config across. Works perfectly, fixing preexisting engraving problem! Disconnected reset button because it was causing crashing. | [Bob](User:parag0n "wikilink") | | 2015-11-23 | Installed red emergency reset button with shielded cable, now works without spurious resets. | [Greg](User:Gregory.morris "wikilink") | | 2015-11-29 | Attempted to fix X-axis issue - Reseated cables, switched stepper drivers, replaced cables entirely, no joy. Can only be motor. | [Bob](User:parag0n "wikilink") | | 2015-12-14 | Replaced X-axis motor, mounted 16T pulley and adjusted tension. Began aligning mirror on motor mount. | [Greg](User:Gregory.morris "wikilink") | | 2015-12-15 | Finished motor mount mirror alignment - still some misalignment on the first mirror. Awaiting 20T pulley. | [Greg](User:Gregory.morris "wikilink") | | 2015-12-27 | Replaced pulley - Working fine. | [Greg](User:Gregory.morris "wikilink") | | 2016-01-07 | Cleaned head mirror and lens. | [Bob](User:parag0n "wikilink") | | 2016-01-11 | Replaced lens as previous one had EXPLODED. | [Bob](User:parag0n "wikilink") | | 2016-02-03 | Cleaned air assist nozzle, lens and mirrors, checked and adjusted alignment. | [Greg](User:Gregory.morris "wikilink") | | 2016-02-27 | Rebuilt cooling system. Now has two flow meters in line, one for the cooling system, and one thats hard-wired into the water detect line on the laser PSU. | [Bob](User:parag0n "wikilink") | | 2016-03-04 | Cleaned air assist nozzle, lens, mirrors, and orange bed . | [Greg](User:Gregory.morris "wikilink") | | 2016-05-26 | Cleaned air assist nozzle, lens, mirrors. Investigated air pump stuck on - not the cable. I suspect a bad relay is the problem. | [Greg](User:Gregory.morris "wikilink") | | 2016-05-29 | Moved air assist pump to relay board channel 3. Now functional. Relay Board Channel 2 should not be used. | [Greg](User:Gregory.morris "wikilink") | | 2016-06-10 | Fixed startup issue with wire loop in connector over lid switch position. User reported issues on Y-axis - motor can be heard to skip steps. This is consistent with the misalignment produced when cutting. Y-axis current and/or acceleration need tweaking. | [Greg](User:Gregory.morris "wikilink") | [Category:Laser Cutter (Orange)](Category:Laser_Cutter_(Orange) "wikilink") [Category:Maintenance Records](Category:Maintenance_Records "wikilink")